# Twin branching in shape memory alloys: a 1D continuum model with energy dissipation effects

### v1

@@ Abstract
We develop a 1D continuum model of twin branching in shape memory alloys. The free energy of the branched microstructure comprises the interfacial and elastic strain energy contributions, both expressed in terms of the average twin spacing treated as a continuous function of the position. The total free energy is then minimized, and the corresponding Euler–Lagrange equation is solved numerically using the finite element method. The model can be considered as a continuum counterpart of the recent discrete model of Seiner et al. (2020), and our results show a very good agreement with that model in the entire range of physically relevant parameters. Furthermore, our continuum setting facilitates incorporation of energy dissipation into the model. The effect of rate-independent dissipation on the evolution of the branched microstructure is thus studied. The results show that significant effects on the microstructure and energy of the system are expected only for relatively small domain sizes.
@@ 1 Introduction
Martensitic transformation, which is the basic deformation mechanism in shape memory alloys, proceeds through formation and evolution of microstructure (Bhattacharya, 2003). Interfaces between the product phase (martensite) and the parent phase (austenite) constitute the most common element of the microstructure. Kinematic compatibility of the two phases at such interfaces is typically achieved through twinning of martensite so that martensite appears in the form of a laminate of two twin-related variants. Twin branching, which is the subject of this work, is a phenomenon frequently observed in experiments (e.g., Chu, 1993; Bronstein et al., 2019; Qin et al., 2023) in which the thickness of the individual twins is gradually reduced towards the austenite–martensite interface.
@@
The theoretical analysis of twin branching started with the seminal contributions of Kohn and Müller (1992, 1994), followed by many related works, mostly in the mathematics community (e.g., Capella and Otto, 2009; Conti, 2000; Chan and Conti, 2015; Dondl et al., 2016; Simon, 2021). This line of research delivers, in particular, the scaling laws for energy and for characteristic dimensions of the branched microstructure. As a follow up, a quantitative model of twin branching in shape memory alloys has been recently developed by Seiner et al. (2020). The model assumes that the branched microstructure consists of a family of self-similar branching generations, as in several earlier models, and refined estimates of the elastic strain energy of branching have been derived for the considered microstructure. The parameters characterizing the actual microstructure are then obtained by minimizing the total free energy. This model will be used as a reference for our model developed below.
@@
In this paper, we propose a new approach to the modeling of twin branching: a continuum 1D model is developed in which the twin spacing is treated as a continuous function of the position. In this continuum setting, the free energy of the system is formulated which includes three contributions: the interfacial energy of twin boundaries, the elastic strain energy of branching, and the energy of elastic strains that accommodate the local incompatibility at the austenite–twinned martensite interface. The total free energy is then minimized, thus delivering the governing equation (ODE) for the twin spacing that is solved numerically using the finite element method. The model can be considered a continuum counterpart of the discrete model of Seiner et al. (2020), and we show that the predictions of the two models agree very well for a wide range of model parameters.
@@
Secondly, for the first time, we consider the energy dissipation associated with the evolution of the branched microstructure. This is facilitated by our continuum formulation, as development of the respective discrete formulation would be difficult, if possible at all. The evolution equation for the twin spacing is obtained by minimization of the rate potential that comprises the rate of the total free energy and the rate-independent dissipation potential. Finite-element computations are then carried out for a high but realistic value of the rate-independent threshold for propagation of twin interfaces. The results suggest that the energy dissipation effects are small or even negligible in most cases. Only for a relatively high value of the dimensionless interfacial energy (which corresponds, for instance, to a small size of the twinned domain) and for a high interface propagation threshold, a significant effect on both the microstructure (twin spacing) and energy can be observed.
@@
The reminder of the paper is organized as follows. In Section 2, the discrete model of twin branching of Seiner et al. (2020) is briefly outlined as a reference for our subsequent developments. In Section 3, the continuum model of twin branching is developed and sample results are presented, including a comparison with the predictions of the model of Seiner et al. (2020). In Section 4, energy dissipation is introduced into the model and the corresponding effects are studied for a few representative cases of evolution of the branched microstructure. Technical details concerning the derivation of the evolution problem and its computational treatment are presented in appendices. A general discussion on the model and its applicability is presented in Section 5, followed by the Conclusion in Section 6.
@@ 2.1 Austenite–twinned martensite interface
In majority of shape memory alloys, the interface between the austenite and the martensite forms such that the martensite is internally twinned, i.e., it has the form of a fine laminate of two twin-related variants of martensite. This is because a single martensite variant is usually not compatible with the austenite and thus cannot form a low-energy (stress-free) austenite–martensite interface (Bhattacharya, 2003). The orientation and other features of the austenite–twinned martensite (A–MM) interface can be predicted by using the crystallographic theory of martensite (Ball and James, 1987; Bhattacharya, 2003), which relies on the kinematic compatibility conditions formulated for the twin interfaces and for the A–MM interface.
@@ 2.2 Discrete model of twin branching
A discrete model of twin branching has been recently developed by Seiner et al. (2020), and this model is briefly summarized below. The model is based on the minimization of the total free energy of the specific branching microstructure shown in Fig. 1(a). It is assumed that the twin spacing is refined through a sequence of self-similar branching segments. Each branching generation, indexed by $i$, is formed by a layer of the corresponding unit cells (segments) that are repeated in a periodic manner with $L_{i}$ denoting the length measured in the direction perpendicular to the A–MM interface and $h_{i}$ denoting the twin spacing measured along the A–MM interface, while the width of the segment is equal to $2h_{i}$, see Fig. 1(b). Here and in the following, we keep the notation used by Seiner et al. (2020), except that the twin spacing is denoted by $h_{i}$ rather than by $d_{i}$. Note that the first layer of the length $L_{0}$ and spacing $h_{0}$ is not branched, see Fig. 1(a).
@@
This result will be used to calibrate our continuum model in Section 3.
@@
3 Continuum model of twin branching
@@ 3.1 Twin spacing as a continuous function
In our continuum model, the twin spacing $h$ is considered to be a continuous function of the position $x$, thus $h=h(x)$ for $0\leq x\leq L$. Considering that real microstructures do not exhibit periodicity nor perfect arrangement of branching generations, as assumed in the discrete model of Seiner et al. (2020), the twin spacing $h(x)$ should be interpreted as the average spacing measured over a representative part of the A–MM interface and determined at a given distance from it. Note that, in accord with Seiner et al. (2020), the twin spacing $h(x)$ is here measured in the direction parallel to the A–MM interface, and the spatial coordinate $x$ is measured in the direction perpendicular to the A–MM interface with $x=0$ at the free boundary and $x=L$ at the A–MM interface, see Fig. 1(a).
The estimate of the elastic strain energy of branching, Eq. (4), will be used to characterize the elastic strain energy in the continuum setting. To this end, we note that the energy density within a single cell depends on the ratio $h_{i}/L_{i}$, see Eq. (6). Now, referring to Fig. 1(c), the spatial derivative of $h(x)$ can be related to the ratio $h_{i}/L_{i}$ through
@@
and this estimate will be used in our continuum model.
@@ 3.2 Governing equations
The governing equations of the continuum model will now be derived through minimization of the total free energy of the system. The free energy density (per unit volume of the branched microstructure) is formulated in the following form,
@@
The first term is the elastic strain energy contribution derived above, see Eq. (9). The second term describes the interfacial energy of twin boundaries and is a continuum counterpart of Eq. (7)2. The free energy function $\phi$ is the density per unit volume, i.e., the density per unit area of the A–MM interface and per unit length along the $x$-axis.
@@
Let us assume that the microstructure of the branched martensite domain is formed such that the total free energy is minimized. This is a common approach in the modeling of martensitic microstructures, followed in particular in the crystallographic theory of martensite (Ball and James, 1987; Bhattacharya, 2003), as well as in the modeling of twin branching (Kohn and Müller, 1992; Seiner et al., 2020). This assumption will be relaxed in Section 4 where the evolution of the microstructure will be considered and the associated energy dissipation will be accounted for. The twin spacing $h(x)$ is thus found by minimizing the total free energy $\Phi$,
@@
where the whole second term (i.e., both $h_{,x}$ and $\delta h$) is evaluated at $x=0$ (this is denoted by the vertical bar with a subscript), and likewise the whole third term is evaluated at $x=L$. Following the standard argument, Eq. (14) implies the following local form of the governing equation,
@@
Note that the boundary condition at $x=0$, Eq. (16)1, is the continuum counterpart of the non-branched first segment in the construction proposed by Seiner et al. (2020), see Fig. 1(a).
@@
The above weak form will be used as a basis for an approximate solution of the problem using the finite element method.
@@ 3.4 Model predictions
The elastic micro-strain energy factor $\Gamma$ is adopted based on the results of the micromechanical analysis carried out using the phase-field method (Tůma and Stupkiewicz, 2016) Out of the four crystallographically distinct microstructures of the A–MM interface, the one considered by Seiner et al. (2020) corresponds to microstructure M3 in the notation used by Tůma and Stupkiewicz (2016). The corresponding values of the energy factor $\Gamma^{\rm e}_{\rm am}$ reported in Fig. 10 in Tůma and Stupkiewicz (2016) range from about 2 MJ/m3 to 5 MJ/m3 depending on the twin spacing and interface thickness parameter of the phase-field model. The lower values of $\Gamma^{\rm e}_{\rm am}$ are associated with a local branching mechanism that becomes effective when the twin spacing is greater than 30 nm, see Fig. 6 in Tůma and Stupkiewicz (2016), and also with a greater interface thickness in the phase-field model. Therefore, we adopt the value of $\Gamma^{\rm e}_{\rm am}\approx 5$ MJ/m3 as a reasonable estimate, which is also consistent with the results of the sharp-interface modeling reported by Stupkiewicz et al. (2007). Note that the factor $\Gamma^{\rm e}_{\rm am}$ in Tůma and Stupkiewicz (2016) corresponds to the twin spacing measured in the direction normal to the twin interfaces, while the factor $\Gamma$ in Eq. (11) corresponds to the twin spacing measured along the A–MM interface. Accordingly, we have $\Gamma=\alpha\Gamma^{\rm e}_{\rm am}$, and thus the value of $\Gamma=3$ MJ/m3 will be used in the following. The corresponding dimensionless parameter is equal to $\bar{\Gamma}=0.345$, see Eq. (19)2.
@@
Figure 2: Twin spacing $h(x)$ for (a) relatively large domain with $L$ varied between 1 mm ($\bar{\gamma}=2\times 10^{-5}$) and 5 mm ($\bar{\gamma}=4\times 10^{-6}$) and (b) relatively small domain with $L$ varied between 4 $\mu$m ($\bar{\gamma}=5\times 10^{-3}$) and 20 $\mu$m ($\bar{\gamma}=10^{-3}$).
Finally, two ranges of the size $L$ of the twinned martensite domain are considered. The first range corresponds to a relatively large domain with $L$ varied between $1$ mm and $5$ mm, where the latter value was used in the main illustrating example of Seiner et al. (2020), see their Fig. 8. The respective results are shown in Fig. 2(a). The second range corresponds to a relatively small domain with $L$ varied between $4\,\mu$m and $20\,\mu$ m, and the respective results are shown in Fig. 2(b). It can be seen that twin branching and the corresponding refinement of the microstructure towards the A–MM interface is more pronounced for larger domains.
Fig. 3 presents a comprehensive comparison of the predictions of the present continuum model with those of the discrete model of Seiner et al. (2020), as reported in their Table 1. The latter results have been reported for the length $L$ varied between $10^{4}$ nm and $10^{7}$ nm (four values) and for the material length $\gamma/\mu$ varied between $10^{-4}$ nm and $10^{-1}$ nm (four values). Apparently, the 16 cases reported by Seiner et al. (2020) collapse with a very good accuracy to seven distinct cases in our dimensionless representation, Eqs. (17), (19) and (20), with the dimensionless twin boundary energy $\bar{\gamma}$ spanning six decades from about $10^{-7}$ to $10^{-1}$.
@@
Figure 3: Predictions of the present continuum model compared to those of the discrete model of Seiner et al. (2020): (a) dimensionless total free energy $\bar{\Phi}$, (b) dimensionless twin spacing $\bar{h}_{0}$ at $x=0$, (c) dimensionless twin spacing $\bar{h}_{L}$ at $x=L$, and (d) the number of branching generations $N$; all as a function of the dimensionless twin boundary energy $\bar{\gamma}$, see Eq. (19)1. The shaded regions indicate the range of values corresponding to $\bar{\Gamma}$ decreased twice (dashed lines) and increased twice (dotted lines).
Fig. 3 shows a very good agreement of the two models over the whole range of $\bar{\gamma}$. In particular, an almost perfect agreement is demonstrated in terms of the dimensionless total free energy $\bar{\Phi}$ and twin spacing $\bar{h}_{0}$ at the free end ($x=0$), see Fig. 3(a,b), except for a small deviation observed for the largest values of $\bar{\gamma}$ (to be commented below). The agreement in terms of $\bar{h}_{L}$, the dimensionless twin spacing at the A–MM interface (at $x=L$), is somewhat worse, see Fig. 3(c). Still it can be considered acceptable, the more that it does not affect the good agreement in the total energy $\bar{\Phi}$ and in the twin spacing $\bar{h}_{0}$.
Fig. 3(d) compares the number of branching generations $N$. In the discrete model of Seiner et al. (2020), $N$ is a positive integer such that $h_{L}=h_{N}=h_{0}/2^{N}$. In the case of our continuum model, $N$ is defined as
@@
and, clearly, takes non-integer values. Consistent with the case of $\bar{h}_{L}$ in Fig. 3(c), a visible difference between the two models is observed also in the case of $N$, however, both models follow the same trend.
For $\bar{\gamma}=10^{-1}$ (the largest value of $\bar{\gamma}$ considered), the continuum model predicts $N=0.27$, with the interpretation that only a fraction of twin lamellae would branch. The discrete model predicts then one branching generation, $N=1$. This may justify the (small) deviation of the two models in terms of $\bar{\Phi}$ and $\bar{h}_{0}$ observed for the largest values of $\bar{\gamma}$, see Fig. 3(a,b). Indeed, when the number of branching generations is small, it is natural to expect that a piecewise constant distribution of twin spacing and a continuous distribution yield somewhat different results. Actually, it may be surprising that the agreement is then still rather good. Note that, adopting the values of parameters $\gamma$, $A$, and $\alpha$ as specified above for the CuAlNi alloy, the case of $\bar{\gamma}=10^{-1}$ would correspond to $L=0.2$$\mu$m, which is a very small and thus non-physical value.
Fig. 3 illustrates also the sensitivity of the model predictions to the value of the elastic micro-strain energy factor $\Gamma$. The shaded regions in each diagram indicate the range of values that correspond to parameter $\bar{\Gamma}$ decreased twice and increased twice with respect to the reference value of $\bar{\Gamma}=0.345$. It can be seen that the total energy $\bar{\Phi}$ and the twin spacing $\bar{h}_{0}$ are practically insensitive to $\bar{\Gamma}$ over a wide range of values of $\bar{\gamma}$, and only for the largest values of $\bar{\gamma}$ they show a moderate sensitivity to $\bar{\Gamma}$. It can also be seen that the agreement with the discrete model is significantly improved for $\bar{\Gamma}=0.690$ (the dotted curves), and even a better, almost perfect agreement would be achieved for $\bar{\Gamma}$ increased further, say, to $\bar{\Gamma}=1$.

### v2

# Twin branching in shape memory alloys: a 1D model with energy dissipation effects

@@ Abstract
We develop a 1D model of twin branching in shape memory alloys. The free energy of the branched microstructure comprises the interfacial and elastic strain energy contributions, both expressed in terms of the average twin spacing treated as a continuous function of the position. The total free energy is then minimized, and the corresponding Euler–Lagrange equation is solved numerically using the finite element method. The model can be considered as a continuous counterpart of the recent discrete model of Seiner et al. (2020), and our results show a very good agreement with that model in the entire range of physically relevant parameters. Furthermore, our continuous setting facilitates incorporation of energy dissipation into the model. The effect of rate-independent dissipation on the evolution of the branched microstructure is thus studied. The results show that significant effects on the microstructure and energy of the system are expected only for relatively small domain sizes.
@@ 1 Introduction
Martensitic transformation, which is the basic deformation mechanism in shape memory alloys, proceeds through formation and evolution of microstructure (Bhattacharya, 2003). Interfaces between the product phase (martensite) and the parent phase (austenite) constitute the most common element of the microstructure. Kinematic compatibility of the two phases at such interfaces is typically achieved through twinning of martensite so that martensite appears in the form of a laminate of two twin-related variants. Twin branching, which is the subject of this work, is a phenomenon frequently observed in experiments (e.g., Chu, 1993; Bronstein et al., 2019; Qin et al., 2023; Zhang et al., 2024) in which the thickness of the individual twins is gradually reduced towards the austenite–martensite interface.
@@
The theoretical analysis of twin branching started with the seminal contributions of Kohn and Müller (1992, 1994), followed by many related works, mostly in the mathematics community (e.g., Conti, 2000; Capella and Otto, 2009; Chan and Conti, 2015; Dondl et al., 2016; Simon, 2021). This line of research delivers, in particular, the scaling laws for energy and for characteristic dimensions of the branched microstructure. As a follow up, a quantitative model of twin branching in shape memory alloys has been recently developed by Seiner et al. (2020). The model assumes that the branched microstructure consists of a family of self-similar branching generations, as in several earlier models, and refined estimates of the elastic strain energy of branching have been derived for the considered microstructure. The parameters characterizing the actual microstructure are then obtained by minimizing the total free energy. This model will be used as a reference for our model developed in this work.
@@
In this paper, we propose a new approach to the modeling of twin branching. A 1D model is developed in which the twin spacing is treated as a continuous function of the position. In this setting, the free energy of the system is formulated which includes three contributions: the interfacial energy of twin boundaries, the elastic strain energy of branching, and the energy of elastic strains that accommodate the local incompatibility at the austenite–twinned martensite interface. The total free energy is then minimized, thus delivering the governing equation (ODE) for the twin spacing that is solved numerically using the finite element method. The model can be considered a continuous counterpart of the discrete model of Seiner et al. (2020), and we show that the predictions of the two models agree very well for a wide range of model parameters.
We call our model a continuous model since the twin spacing is treated as a continuous function of the position and the governing equation can be formulated as an ODE. This is in contrast to the model of Seiner et al. (2020), which we call a discrete model, since that model considers a discrete number of branching generations thus leading to a step-like dependence of the twin spacing on the position. Moreover, the solution is obtained from a discrete minimization problem, in particular, involving minimization with respect to the (integer) number of branching generations.
A model similar to our model has been recently [^1] The paper of Zhang et al. (2024) was published after our paper was submitted for publication, and we were not aware of this work while preparing the first version of our paper. We have added the reference to this work at the revision stage. presented by Zhang et al. (2024) for the analysis of experimental data on austenite–martensite interfacial patterns in NiMnGa reported in the reference. That model also considers the twin spacing as a continuous function of the position and adopts the same form of the free energy density in the branched martensite. The difference is that our model includes the energy of elastic micro-strains at the A–MM interface, which is disregarded in the model of Zhang et al. (2024). This leads to different boundary conditions for the corresponding ODE. The model of Zhang et al. (2024), once fitted to their experimental data, correctly describes the profile of the twin spacing in the vicinity of the austenite–martensite interface.
@@
Secondly, for the first time, we consider the energy dissipation associated with the evolution of the branched microstructure. This is largely facilitated by our continuous formulation (note that development of the respective discrete formulation of the evolution problem may not be readily feasible). The evolution equation for the twin spacing is obtained by minimization of the rate potential that comprises the rate of the total free energy and the rate-independent dissipation potential. Finite-element computations are then carried out for a high but realistic value of the rate-independent threshold for propagation of twin interfaces. The results suggest that the energy dissipation effects are small or even negligible in most cases. Only for a relatively high value of the dimensionless interfacial energy (which corresponds, for instance, to a small size of the twinned domain) and for a high interface propagation threshold, a significant effect on both the microstructure (twin spacing) and energy can be observed.
@@
The remainder of the paper is organized as follows. In Section 2, the discrete model of twin branching of Seiner et al. (2020) is briefly outlined as a reference for our subsequent developments. In Section 3, the continuous model of twin branching is developed and sample results of finite-element computations are presented, including a comparison with the predictions of the model of Seiner et al. (2020). As a side note, an analytical solution to this (static) problem is discussed in A. In Section 4, energy dissipation is introduced into the model and the corresponding effects are studied for a few representative cases of evolution of the branched microstructure. Technical details concerning the derivation of the evolution problem and its computational treatment are presented in B and C, respectively. A general discussion on the model and its applicability is presented in Section 5, followed by concluding remarks in Section 6.
@@ 2.1 Austenite–twinned martensite interface
In majority of shape memory alloys, the interface between the austenite and the martensite forms such that the martensite is internally twinned, i.e., it has the form of a fine laminate of two twin-related variants of martensite. This is because a single martensite variant is usually not compatible with the austenite and thus cannot form a low-energy (stress-free) austenite–martensite interface (Bhattacharya, 2003). The orientation and other features of the austenite–twinned martensite (A–MM) interface can be predicted by using the crystallographic theory of martensite (Wechsler, 1953; Ball and James, 1987; Bhattacharya, 2003), which relies on the kinematic compatibility conditions formulated for the twin interfaces and for the A–MM interface.
@@ 2.2 Discrete model of twin branching
A discrete model of twin branching has been recently developed by Seiner et al. (2020), and this model is briefly summarized below. The model is based on the minimization of the total free energy of the specific branching microstructure shown in Fig. 1(a). It is assumed that the twin spacing is refined through a sequence of self-similar branching segments. Each branching generation, indexed by $i$, is formed by a layer of the corresponding unit cells (segments) that are repeated in a periodic manner with $L_{i}$ denoting the length measured in the direction perpendicular to the A–MM interface and $h_{i}$ denoting the twin spacing measured along the A–MM interface, while the width of the segment is equal to $2h_{i}$, see Fig. 1(b). Here and in the following, we keep the notation used by Seiner et al. (2020), except that the twin spacing is denoted by $h_{i}$ rather than by $d_{i}$ and $\Gamma$ replaces $G_{AB}$. Note that the first layer of the length $L_{0}$ and spacing $h_{0}$ is not branched, see Fig. 1(a).
@@
This result will be used to calibrate our continuous model in Section 3.
@@
3 Continuous model of twin branching
@@ 3.1 Twin spacing as a continuous function
In our model, the twin spacing $h$ is considered to be a continuous function of the position $x$, thus $h=h(x)$ for $0\leq x\leq L$. Considering that real microstructures do not exhibit periodicity nor perfect arrangement of branching generations, as assumed in the discrete model of Seiner et al. (2020), the twin spacing $h(x)$ should be interpreted as the average spacing measured over a representative part of the A–MM interface and determined at a given distance from it. Note that, in accord with Seiner et al. (2020), the twin spacing $h(x)$ is here measured in the direction parallel to the A–MM interface. The spatial coordinate $x$ is measured in the direction perpendicular to the A–MM interface with $x=0$ at the free boundary and $x=L$ at the A–MM interface, see Fig. 1(a).
The estimate of the elastic strain energy of branching, Eq. (4), will be used to characterize the elastic strain energy in the continuous setting. To this end, we note that the energy density within a single cell depends on the ratio $h_{i}/L_{i}$, see Eq. (6). Now, referring to Fig. 1(c), the spatial derivative of $h(x)$ can be related to the ratio $h_{i}/L_{i}$ through
@@
and this estimate will be used in our model.
@@ 3.2 Governing equations
The governing equations of the model will now be derived through minimization of the total free energy of the system. The free energy density (per unit volume of the branched microstructure) is formulated in the following form,
@@
The first term is the elastic strain energy contribution derived above, see Eq. (9). The second term describes the interfacial energy of twin boundaries and is a continuous counterpart of Eq. (7)2. The free energy function $\phi$ is the density per unit volume, i.e., the density per unit area of the A–MM interface and per unit length along the $x$-axis.
@@
We assume that the microstructure of the branched martensite domain is formed such that the total free energy is minimized. This is a common approach in the modeling of martensitic microstructures, followed in particular in the crystallographic theory of martensite (Ball and James, 1987; Bhattacharya, 2003), as well as in the modeling of twin branching (Kohn and Müller, 1992; Seiner et al., 2020). This assumption will be relaxed in Section 4 where the evolution of the microstructure will be considered and the associated energy dissipation will be accounted for. The twin spacing $h(x)$ is thus found by minimizing the total free energy $\Phi$,
@@
where the whole second term is evaluated at $x=0$ (i.e., both $h_{,x}$ and $\delta h$, as denoted by the vertical bar with a subscript), and likewise the whole third term is evaluated at $x=L$. Following the standard argument, Eq. (14) implies the following local form of the governing equation,
@@
Note that the boundary condition at $x=0$, Eq. (16)1, is a continuous counterpart of the non-branched first segment in the construction proposed by Seiner et al. (2020), see Fig. 1(a).
Let us multiply Eq. (15) by $h_{,x}$ (and by $A$) and integrate it once over $x$. As a result, the following property of the solution of the problem at hand is obtained,
$\frac{1}{2}A(h_{,x})^{2}-\frac{2\gamma}{\alpha h}=C,$ (17)
where $C$ is an integration constant. This is the property of equipartition of energy (Kohn and Müller, 1992), which shows that the difference between the elastic strain energy $\frac{1}{2}A(h_{,x})^{2}$ and the interfacial energy $2\gamma/(\alpha h)$ is independent of $x$.
@@
The above weak form will be used as a basis for an approximate solution of the problem using the finite element method. Note that the problem at hand admits an analytical solution that is derived in A. However, the solution requires solving two nonlinear equations numerically. Since the finite element method is indispensable for the evolution problem studied in Section 4, it is also used for this static case.
@@ 3.4 Model predictions
The last model parameter to be determined is the elastic micro-strain energy factor $\Gamma$. In their supplementary material, Seiner et al. (2020) provide an upper-bound estimate of the elastic energy of closure domains that could be used to compute $\Gamma$. Instead of following this route, we identify $\Gamma$ from the results reported in their Fig. 8(a). Specifically, for $N=0$, i.e., for the case with no branching, the total energy in the twinned martensite, denote it by $\Phi^{\ast}$ ($E^{(0)}$ in their notation), has been reported to be equal to 270 N/m. On the other hand, for a constant $h$, the total energy is given by $\Phi^{\ast}=2\gamma L/(\alpha h)+\Gamma h$, which upon minimization with respect to $h$ is found equal to $\Phi^{\ast}=\sqrt{8L\Gamma\gamma/\alpha}$. Finally, parameter $\Gamma=10.3$ MJ/m3 can be computed from the latter formula. The corresponding dimensionless parameter is equal to $\bar{\Gamma}=1.2$, see Eq. (20)2.
As a reference, another value of parameter $\Gamma$ will also be used in the computations. This will, in particular, illustrate the sensitivity of the results to the value of $\Gamma$. To this end, parameter $\Gamma$ will be estimated based on the results of the micromechanical analysis carried out using the phase-field method (Tůma and Stupkiewicz, 2016). Out of the four crystallographically distinct microstructures of the A–MM interface, the one considered by Seiner et al. (2020) corresponds to microstructure M3 in the notation used by Tůma and Stupkiewicz (2016). The corresponding values of the energy factor $\Gamma^{\rm e}_{\rm am}$ reported in Fig. 10 in Tůma and Stupkiewicz (2016) range from about 2 MJ/m3 to 5 MJ/m3 depending on the twin spacing and interface thickness parameter of the phase-field model. The lower values of $\Gamma^{\rm e}_{\rm am}$ are associated with a local branching mechanism that becomes effective when the twin spacing is greater than 30 nm, see Fig. 6 in Tůma and Stupkiewicz (2016), and also with a greater interface thickness in the phase-field model. Therefore, we adopt the value of $\Gamma^{\rm e}_{\rm am}\approx 5$ MJ/m3 as a reasonable estimate, which is also consistent with the results of the sharp-interface modeling reported by Stupkiewicz et al. (2007). Note that the factor $\Gamma^{\rm e}_{\rm am}$ in Tůma and Stupkiewicz (2016) corresponds to the twin spacing measured in the direction normal to the twin interfaces, while the factor $\Gamma$ in Eq. (11) corresponds to the twin spacing measured along the A–MM interface. Accordingly, we have $\Gamma=\alpha\Gamma^{\rm e}_{\rm am}$, and thus the value of $\Gamma=3$ MJ/m3 will be used in the following (the corresponding dimensionless parameter is equal to $\bar{\Gamma}=0.35$). This value is three times lower than that used by Seiner et al. (2020), and we believe it is a more accurate estimate of $\Gamma$ than their upper-bound estimate, as it is determined by a kind of shape optimization delivered by the phase-field method and it satisfies the continuum equilibrium equations, in contrast to the upper-bound estimate of Seiner et al. (2020).
Finally, two ranges of the size $L$ of the twinned martensite domain are considered. The first range corresponds to a relatively large domain with $L$ varied between $1$ mm and $5$ mm, where the latter value was used in the main illustrating example of Seiner et al. (2020), see their Fig. 8. The respective results are shown in Fig. 2(a). The second range corresponds to a relatively small domain with $L$ varied between $4\,\mu$m and $20\,\mu$m, and the respective results are shown in Fig. 2(b). It can be seen that twin branching and the corresponding refinement of the microstructure towards the A–MM interface is more pronounced for larger domains. The effect of parameter $\bar{\Gamma}$ characterizing the elastic micro-strain energy at the A–MM interface is negligible in the case of relatively large domain, as shown in Fig. 2(a), except in the vicinity of the A–MM interface, however, the corresponding difference is not discernible in Fig. 2(a). In the case of the relatively small domain, Fig. 2(b), the effect of $\bar{\Gamma}$ is visible, and it is pronounced more at the A–MM interface, at $x=L$, than at the free end, at $x=0$.
@@
Figure 2: Twin spacing $h(x)$ for (a) relatively large domain with $L$ varied between 1 mm ($\bar{\gamma}=2\times 10^{-5}$) and 5 mm ($\bar{\gamma}=4\times 10^{-6}$) and (b) relatively small domain with $L$ varied between 4 $\mu$m ($\bar{\gamma}=5\times 10^{-3}$) and 20 $\mu$m ($\bar{\gamma}=10^{-3}$). In panel (a), the solid lines ($\bar{\Gamma}=1.2$) and the dashed lines ($\bar{\Gamma}=0.35$) overlap.
Fig. 3 presents a comprehensive comparison of the predictions of the present continuous model with those of the discrete model of Seiner et al. (2020), as reported in their Table 1. Their results have been reported for the length $L$ varied between $10^{4}$ nm and $10^{7}$ nm (four values) and for the material length $\gamma/\mu$ varied between $10^{-4}$ nm and $10^{-1}$ nm (four values). Apparently, the 16 cases reported by Seiner et al. (2020) collapse with a very good accuracy to seven distinct cases in our dimensionless representation, Eqs. (18), (20) and (21), with the dimensionless twin boundary energy $\bar{\gamma}$ spanning six decades from about $10^{-7}$ to $10^{-1}$.
@@
Figure 3: Predictions of the present continuous model compared to those of the discrete model of Seiner et al. (2020): (a) dimensionless total free energy $\bar{\Phi}$, (b) dimensionless twin spacing $\bar{h}_{0}$ at $x=0$, (c) dimensionless twin spacing $\bar{h}_{L}$ at $x=L$, and (d) the number of branching generations $N$; all as a function of the dimensionless twin boundary energy $\bar{\gamma}$, see Eq. (20)1.
Fig. 3 shows the dimensionless total free energy $\bar{\Phi}$ as a function of $\bar{\gamma}$ and the dimensionless twin spacing at the free end ($x=0$) and at the A–MM interface ($x=L$), denoted by $\bar{h}_{0}$ and $\bar{h}_{L}$, respectively. Moreover, the number of branching generations $N$ is shown in Fig. 3(d). In the discrete model of Seiner et al. (2020), $N$ is a positive integer such that $h_{L}=h_{N}=h_{0}/2^{N}$. In the case of our continuous model, an analogous quantity $N$ can be defined,
@@
which, clearly, takes non-integer values. Since branching is assumed to proceed in a continuous manner, rather than in the form of distinct branching generations, this number is meant to quantify the microstructure refinement, and does not indicate the actual number of distinct branching generations.
As discussed above, the computations have been carried out for two values of parameter $\bar{\Gamma}$. An almost perfect agreement with the results of Seiner et al. (2020) is obtained in the entire range of $\bar{\gamma}$ when $\bar{\Gamma}=1.2$, i.e., when the same input parameters are used in the two models, see Fig. 3. This is a remarkable result, in particular, for the largest values of $\bar{\gamma}$ for which only one branching generation is predicted by the model of Seiner et al. (2020), and the major characteristics of the solution are correctly represented by our continuous model.
The results corresponding to a lower value of $\bar{\Gamma}=0.35$ are also included in Fig. 3, see the dashed lines, and illustrate the sensitivity of the results to $\bar{\Gamma}$. It follows that the total energy $\bar{\Phi}$ and the twin spacing $\bar{h}_{0}$ at the free end are only weakly affected by $\bar{\Gamma}$ for the largest values of $\bar{\gamma}$ and are practically insensitive to $\bar{\Gamma}$ for $\bar{\gamma}$ varied between $10^{-7}$ and $10^{-3}$. The twin spacing at the A–MM interface, $\bar{h}_{L}$, is visibly affected by $\bar{\Gamma}$, and this is expected since parameter $\bar{\Gamma}$ governs the elastic micro-strain energy at the A–MM interface and hence directly impacts $\bar{h}_{L}$. As a result, also the number of twin branching generations $N$ is visibly affected by $\bar{\Gamma}$. For $\bar{\gamma}=10^{-1}$ (the largest value of $\bar{\gamma}$ considered), the continuous model predicts $N=0.27$ with the interpretation that only a fraction of twin lamellae would branch (the discrete model predicts then one branching generation, $N=1$).
The theory of Kohn and Müller (1992) predicts that, in the case of the self-similar construction, i.e., for small $\gamma$ and large $\Gamma$, the total free energy $\Phi$ and the twin spacing at the free end, $h_{0}$, scale according to $\Phi\sim\mu^{1/3}\gamma^{2/3}L^{1/3}$ and $h_{0}\sim\mu^{-1/3}\gamma^{1/3}L^{2/3}$, respectively. Seiner et al. (2020) have checked that their results also satisfy these scaling laws with a good accuracy, see their Fig. 8(d,e). Now, in our dimensionless representation, the scaling laws of Kohn and Müller (1992) take a particularly simple form, namely
$\bar{\Phi}\sim\bar{\gamma}^{2/3},\qquad\bar{h}_{0}\sim\bar{\gamma}^{1/3}.$ (24)
As can be seen from Fig. 3(a,b), the results of Seiner et al. (2020) indeed follow these two scaling laws. Apparently, it is also the case of our results for $\bar{\Gamma}=1.2$. However, for $\bar{\Gamma}=0.35$, our results deviate from the two scaling laws for $\bar{\gamma}>10^{-3}$, thus indicating a deviation from self-similar regime of Kohn and Müller (1992).